OpenStudy (anonymous):

If x varies inversely as y and directly as t, and x = 12 when t = 10 and y = 25, find y when x is 6 and t = 3.

OpenStudy (shubhamsrg):

given x=c t/y for some constant c also 12 = c 10/25 you need to solve for y from : 6= c 3/y easier now?

OpenStudy (anonymous):

no its not an answer choice

OpenStudy (anonymous):

given x=c t/y for some constant c also 12 = c 10/25 you need to solve this for c : c=25*12/10=30 so the equation is: x=30 t/y now substitute the values of x and t to find corresponding y: 6=30*3/y y=90/6=15

OpenStudy (shubhamsrg):

well even if you didnt want to solve for c,, you can simply divide the 2 eqns..

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

thanks myko

OpenStudy (anonymous):

yw